An investment will pay $100 at the end of each of the next 3 years, $200 at the end of Year 4, $300 at the end of Year 5, and $500 at the end of Year 6. If other investments of equal risk earn 9% annually, what is this investment's present value? Its future value? Do not round intermediate calculations. Round your answers to the nearest cent.

Present Value ?

Future Value?
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=100/1.09+100/1.09^2+100/1.09^3+200/1.09^4+300/1.09^5+500/1.09^6

=$887.93(Approx).

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

Future value=887.93*(1.09)^6

=$1489.14(Approx).